### v1

# IAFormer: Interaction-Aware Transformer network for collider data analysis

Abstract
In this paper, we introduce IAFormer, a novel Transformer-based architecture that efficiently integrates pairwise particle interactions through a dynamic sparse attention mechanism. The IAformer has two new mechanisms within the model. First, the attention matrix depends on predefined boost invariant pairwise quantities, reducing the network parameter significantly from the original particle transformer models. Second, IAformer incorporate the sparse attention mechanism by utilizing the “differential attention”, so that it can dynamically prioritizes relevant particle tokens while reducing computational overhead associated with less informative ones. This approach significantly lowers the model complexity without compromising performance. Despite being computationally efficient by more than an order of magnitude than the Particle Transformer network, IAFormer achieves state-of-the-art performance in classification tasks on the Top and quark-gluon datasets. Furthermore, we employ AI interpretability techniques, verifying that the model effectively captures physically meaningful information layer by layer through its sparse attention mechanism, building an efficient network output that is resistant to statistical fluctuations. IAformer highlights the need to sparse attention in any Transformer analysis to reduce the network size while improving its performance.
1 Introduction
Analyzing LHC collision data to uncover intriguing physics phenomena, selecting an appropriate jet reconstruction approach and tagging algorithm is essential, as they must align with the expected characteristics of the final state. For scenarios involving the production of a high momentum, hadronically decaying heavy particle such as a W, Z, or a top quark, a large radius jet is often used to capture entire particles arising from the heavy particle decay.
The recognition of fat jets with the multi-prong structure was first noted in Ref. [1], where jet substructure techniques emerged as a powerful method for suppressing QCD backgrounds, thereby advancing Beyond Standard Model (BSM) searches at the LHC. The pronged structure associated with different heavy resonant particles is dictated by the final state products of their decays. Over time, jet substructure methodologies have evolved, leading to Deep Learning (DL) based classification techniques for tagging heavy particles [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. These approaches are now widely applied in LHC analyses conducted by ATLAS and CMS [22, 23, 24, 25, 26, 27]. They are also employed in scenarios with multiple heavy particle final states [28, 29, 30] and in extracting fundamental parameters of the theory [31, 32, 33, 34, 35]. In contrast, processes that generate spatially isolated quarks or gluons are typically reconstructed using one or more small radius jets. This is followed by flavor tagging, a technique that identifies the type of initiating parton, whether an up, down, strange, charm, or bottom quark, or a gluon. Bottom and charm hadrons (primarily B and D hadrons) have relatively long lifetimes, allowing them to travel a short distance within the detector before decaying. Their decays produce displaced tracks originating from a secondary vertex, distinct from the primary collision point. By leveraging machine-learning techniques, this vertex displacement information, combined with subsequent decay properties, significantly improves tagging efficiency [36, 37, 38, 39]. Additionally, studies have explored the potential for strange-quark identification using ML techniques [40, 41, 42]. Since gluons belong to the octet representation of QCD SU(3) symmetry, their jet characteristics differ from those of quarks, which are in the triplet representation. Consequently, ML-based methods have also been developed to distinguish gluon-initiated jets from quark-initiated ones [43, 44, 45, 46, 47, 48].
Over the past few decades, significant advancements have been made in jet identification techniques at hadron colliders. In particular, DL methods have transformed how we analyze and utilize jet-related observables. Treating a jet as a single entity provides limited information about its origin, whereas a more detailed analysis of its internal structure and constituent particles offers valuable insights into the nature of the initiating particle. This has led to the development of various techniques collectively referred to as jet tagging, which aim to classify jets based on their underlying physics.
Recently, Transformer networks have demonstrated significant improvements in jet tagging tasks [49, 50, 21]. In their standard architecture, Transformers process jet constituents along with their corresponding features, which include kinematic information for each individual hadron inside the jet. In this setup, the network assigns weights to each particle based on its feature importance in the overall classification decision. However, incorporating features derived from particle-pair interactions can further enhance network performance, as these features play a crucial role in capturing the radiation pattern inside the jet [10]. Efforts to integrate pairwise interactions into the attention mechanism have involved either using an interaction matrix as a bias for the computed attention scores [49] or replacing the Query and Key matrices with the interaction matrix itself [50]. In both approaches, the number of attention heads is constrained by the number of features in the interaction matrix. Additionally, the interaction matrix remains fixed across attention heads and is injected into each attention layer without updates. This rigid structure can disrupt the learning of attention patterns across different layers, potentially limiting the network overall effectiveness. IAFormer addresses these challenges by allowing the interaction matrix to be updated across different attention heads and Transformer layers. Furthermore, the interaction matrix is refined through skip connections and propagated to later attention layers.
Another key innovation in IAFormer lies in its ability to dynamically focus on the most important particles while suppressing attention to less relevant ones. This is achieved through a dynamic sparse attention mechanism, implemented as the difference between two copies of the learned interaction matrix [51]. By encouraging the network to learn distinct representations for these copies, IAFormer prioritizes essential hadrons while de-emphasizing soft radiation, which often exhibits similar patterns in both signal and background.
The IAFormer architecture demonstrates superior performance compared to previously introduced Transformer-based models when evaluated on top tagging and quark-gluon tagging tasks, achieving this with a significantly reduced network size. Although the network is validated against state-of-the-art jet tagging models, IAFormer is designed as a general framework that can be adapted for various classification tasks. The implementation is publicly available at IAFormer GitHub, and detailed usage instructions are provided in the appendix.
This paper is organized as follows; Section 2 discusses various Transformer network architectures, including the structure of IAFormer. Section 3 presents validation results, comparing IAFormer with existing networks for top tagging and quark-gluon tagging tasks. In Section 4, we employ different interpretability methods to analyze the learned attention patterns in IAFormer during the top tagging task. Section 5 provides a discussion and conclusion.
2 Transformers for jet tagging
Transformers were originally introduced for natural language processing, particularly for sequence-to-sequence tasks like machine translation [52], which were later modified for collider physics. At the core of the Transformer is the self-attention mechanism, which dynamically assigns importance to different input components. One of the most significant advantages of self-attention is its ability to capture long-range dependencies, as it considers all elements of the input simultaneously rather than relying on sequential processing. Moreover, modifications to the self-attention formula by incorporating the particle pair interactions enhance the network’s performance.
In jet tagging tasks, the outcome of an event should remain unchanged regardless of the order in which particles are processed. Therefore, the designed DL networks should treat particles as an unordered set, commonly referred to as a “particle cloud”. By representing hardons inside the jet in a permutation-invariant manner, particle cloud models overcome the challenges posed by combinatorial ambiguities. The particle-based methods also provide the flexibility to incorporate additional particle attributes, such as particle identification and vertex information.
Several machine learning architectures have been developed for jet tagging under the permutation invariant framework, including Deep Sets [53], Edge Convolution Networks [54], and Transformers [49, 29, 55, 56, 16, 20]. Deep Sets can achieve state-of-the-art performance but require a large latent space, making them computationally demanding. Edge Convolution Neural Networks (EDGCNN) address this challenge by leveraging local neighborhood information within the particle cloud to enhance learning. For the case of particle transformers, attention weights are computed for every particle interaction within the dataset and outputs are aggregated to keep the permutation information, so that the network output remains unaffected by the order of the input. However, self-attention comes with computational challenges, particularly its quadratic scaling with input size. Various improvements, such as sparse attention [57], have been proposed to address this issue.
In the following, we first introduce the Plain Transformer and its modified version including the interaction matrix, then present the new network, IAFormer. IAFormer significantly reduces the computational burden of attention matrix calculation by incorporating prefixed pairwise information with trainable weights. Moreover, it reduces statistical fluctuations by introducing a dynamic spatial attention mechanism, referred to as differential attention.
2.1 Plain Transformer
The core component of a transformer model is the attention matrix, which is a weighted sum of the input elements. The weights are determined based on the relevance of each element to the others. By ensuring that the attention weights are computed in a permutation-invariant manner, the resulting model is able to effectively process sets.
Mathematically, the attention mechanism operates as follows: The input to the multi-head attention $X$ has an unordered grid-like structure with dimensions, $I\times J$, with $I$ indicating the number of particles and $J$ indicating the relevant features for each particle in the dataset.
For given $X_{ij}$, the feature vector $X_{*j}$ is embedded to a higher dimension via an MLP. The embedding to a higher dimension helps the network to capture meaningful complex patterns and nonlinear relationships that may not be apparent in the original feature space; therefore, it enhances the model’s ability to distinguish different classes for deep learning tasks like classification and anomaly detection. High-dimensional embeddings also provide richer representations that facilitate downstream tasks such as contrastive learning and clustering. The size of the embedded features is determined by the number of neurons in the final MLP layer as
$X^{\rm embed}_{i,j}=\sigma\left(X^{(l-1)}_{i,j^{\prime}}W^{l}_{j^{\prime},j}+b^{l}_{i}\right)\,,$ (1)
with $l$ is the number of the MLP layer, $\sigma$ is a nonlinear activation function, and $W$ and $b$ are the layer weights and bias.
Once the input features are mapped to a higher dimension, the inputs are passed to three separate linear layers forming the Query, Key, and Value matrices as
$Q^{I\times I}_{i,j}\equiv X^{\rm embed}_{i,j^{\prime}}\cdot W^{Q}_{j^{\prime},j}\,,\hskip 11.38109ptK^{I\times J}_{i,j}\equiv X^{\rm embed}_{i,j^{\prime}}\cdot W^{K}_{j^{\prime},j}\,,\hskip 11.38109ptV^{I\times J}_{i,j}\equiv X^{\rm embed}_{i,j^{\prime}}\cdot W^{V}_{j^{\prime},j}\,,$ (2)
where $Q$, $K$ and $V$ are the query, key, and value matrices, respectively, used to compute the attention for the dataset. The suffix $I\times J$ indicates the dimension of the matrices, where $I$ is the number of particle tokens and $J$ is the number of features. The weight matrices have the dimensions of $I\times I$ to preserve the dimensions of the embedded input dataset. The attention score $\alpha$ is defined as:
$\alpha^{I\times I}\equiv\text{softmax}\left(\frac{Q\cdot K^{T}}{\sqrt{d}}\right)=\frac{\text{exp}(Q\cdot(K)^{T}/\sqrt{d})}{\sum\text{exp}(Q\cdot(K)^{T}/\sqrt{d})}\,$ (3)
where in the last line sum runs over all components $I\times I$, suffix $I\times I$ of $\alpha$ indicates the dimension of the attention weight matrix $\alpha$, and $d$ is the length of the particle tokens.
The attention score represents the relative importance of one element in a sequence with respect to another element in the context of a given task. The definition of $\alpha$ with softmax on the scaled dot product ensures that the attention distributed across different elements sums to one. This self-attention mechanism determines how much focus should be given to different input elements when making predictions. Unlike traditional sequence models, which rely on predefined relations among the inputs, attention mechanisms dynamically assign different weights to different parts of the input, allowing the model to emphasize the most relevant features without prejudice.
The attention output is computed as
$\mathcal{Z}_{i,j}=\alpha_{i,i^{\prime}}\cdot V_{i^{\prime},j}\,.$ (4)
The attention output matrix has the same dimensions as the first input dataset, $X^{I\times J}$. Each transformer layer incorporates a multi-head attention mechanism, which combines attention heads to enable parallel and multi-dimensional processing of inputs. This approach enhances the network performance as multi-head attention increases the expressiveness of the model. Each head learns its own projection of the input data, which means the model can explore different subspaces of the feature space. This is particularly useful for capturing both local and global dependencies within the input. As the input dataset is structured in an unordered way, for example, some heads may focus on short-range dependencies, with low $\Delta R$, while others may capture long-range relationships, with larger $\Delta R$.
The outputs from the different attention heads are linearly combined to produce the final output of the multi-head attention layer. The output of the multi-head cross-attention is given by
$\mathcal{O}^{I\times J}=\text{concat}\left(\mathcal{Z}_{I\times J}^{1},\mathcal{Z}_{I\times J}^{2},\ldots,\mathcal{Z}_{I\times J}^{n}\right)\cdot W_{n\cdot J\times J}\,,$ (5)
with $W_{n\cdot J\times J}$ is the learnable linear transformation matrix, so that the final output $\mathcal{O}$ maintains the dimensions of the input dataset. When the outputs of different attention heads are concatenated and projected back into the model feature space, the model effectively combines multiple views of the data into a unified representation. This fusion of different perspectives contributes to the robustness and flexibility of Transformer models.
Finally, the attention output is used to scale the input dataset via a skip connection
$\widetilde{X}^{I\times J}=X^{I\times J}+\mathcal{O}^{I\times J}\,.$ (6)
The transformed dataset, $\widetilde{X}$, represents the relative importance of each element within the dataset, captured from all particles, to the network decision making. Moreover, the transformed dataset has the same dimensions as the input one, which enables stacking more Transformer layers for better performance.
2.2 Transformer with interaction matrix
[…]
$\alpha^{I\times I}=\text{softmax}\left(\frac{Q^{I\times J}\cdot(K^{I\times J})^{T}}{\sqrt{d}}+{\mathcal{I}}^{I\times I}\right)\,,$ (7)
with ${\mathcal{I}}_{i,i}$ is the pairwise interaction matrix for a single head after the embedding. In this structure, the interaction matrix acts as a bias in the attention score. While this approach effectively integrates the interaction matrix into the attention score, it has certain limitations. First, the number of attention heads has to be increased with the number of features of the interaction matrix. If the interaction matrix has a large number of features, the number of attention heads increases accordingly, leading to a significant rise in network parameters. This occurs because $Q$ and $K$ must accommodate the same number of attention heads in this setup. Second, all Transformer layers use the same fixed interaction matrix as a bias, preventing the model from learning updated representations in deeper layers, leading to lower performance.
Another approach is to replace $Q$ and $K$ matrices by the interaction matrix [50]. This modifies the attention score as
$\alpha^{I\times I}=\text{softmax}\left({\mathcal{I}}^{I\times I}\right)\,.$ (8)
Omitting the Query and Key matrices eliminates the necessity of expanding the input feature dimensions of $Q\cdot K^{T}$. Therefore, this structure mitigates the issue of increasing network parameters as the number of features in the pairwise interaction matrix grows.
While this approach reduces network complexity and improves performance, it has a key limitation. Since the interaction matrix is independently added to each Transformer layer. This leads to the mismatch in feature representations between the pairwise interaction matrix and the input $X_{i,j}$. As the attention mechanism assigns weights to all particle tokens in the dataset, this misalignment can lead to assigning attention to irrelevant particles, such as soft hadrons.
2.3 IAFormer
IAFormer is a transformer-based architecture, but significant modifications have been made compared to the ParT setup, as illustrated in Figure 2.3.2. Instead of computing the attention score through the multiplication of the Query and Key matrices, IAFormer applies the softmax function directly to a trainable interaction matrix of the form $W{\mathcal{I}}$. This interaction matrix is independently optimized for each attention head within the individual attention layers, enabling the network to dynamically learn discriminative patterns among different jets with effectively reduced size.
Figure 1: Schematic architecture of IAFormer network.
Additionally, a dynamic sparse attention pattern is included following the idea of “differential attention” [51]. The differential attention utilises the attention scores as the difference between two separate softmax attention maps. The subtraction cancels noise, promoting the sparse attention patterns dynamically. The model has a trainable parameter $\beta$ to control the level of the subtraction. The parameter is initialized randomly and shared across all attention heads within each attention layer. This mechanism suppresses attention scores for less relevant particle tokens, enhancing the model’s focus on important interactions.
2.3.1 Dynamic sparse attention via differential attention
Sparse attention is a technique used in Transformer models to reduce the computational complexity of the self-attention mechanism by limiting the number of key-query pairs attend to the classification. Instead of computing attention over all tokens in a sequence, sparse attention selectively attends to a subset of tokens based on predefined patterns, learned structures, or efficient approximations. This significantly lowers the memory and computational cost from quadratic to linear or sub-quadratic complexity, making it possible to process longer sequences efficiently.
Different types of sparse attention have been introduced such as fixed pattern and learnable sparse attention. Fixed pattern sparse attention is a type of sparse attention mechanism where each token attends only to a predefined subset of tokens based on a fixed rule.[58, 59, 60]. On the other hand, Learnable sparse attention dynamically determines which tokens to attend to based on learned patterns. Learnable sparse attention allows the model to adaptively focus on the most relevant tokens in different contexts, leading to improved efficiency. In this paper we utilize a dynamic sparse attention mechanism incorporated in transformer called “differential attention” which was first introduced in [51].
$\alpha_{i,i^{\prime}}=\rm softmax(W_{1}\cdot\mathcal{I}_{i,j})-\beta\ \rm softmax(W_{2}\cdot\mathcal{I}_{i,j})\,,$ (9)
with $\beta$ is a learnable vector and $W_{1},W_{2}$ are two learnable matrices. During the training process, the value of $\beta$ vector is optimized to demote the attention score assigned to less relevant tokens, resulting in the attention score matrix to be sparse. This results in an implicit sparsity, where tokens gradually focus only on meaningful interactions while ignoring less relevant ones. The self-attention mechanism is guided by this dynamic sparsity, reducing computational overhead while preserving essential long-range dependencies. The IAFormer shows better performance compared with ParT with notably smaller number of parameters.
Despite the advantages of dynamic sparse attention, a challenge remains in controlling the value of the $\beta$ parameter. In the current setup, we clip $\beta$ to lie within the range [0, 1]. Additionally, the attention score $\alpha_{i,j}$ can sometimes take negative values. In such cases, it becomes unclear how to interpret it as an attention probability.
2.3.2 Network structure and implementation
The IAFormer architecture comprises a data embedding block, attention-based layers, and a final MLP for classification as can be seen in Figure . Unlike original Transformer-based models, our approach does not require a class token to aggregate learned information across the layers. Instead, we replace it with a simple pooling operation applied to the final layer output.  [^1] The class token is particularly important in vision tasks, where an input image is divided into small patches, as it helps capture the global structure by aggregating information from Transformer layers. However, since our dataset consists of particle clouds, we replace the class token with average pooling [61]. Two input datasets are used, particle kinematic with dimension $(100,11)$ and interaction matrix with dimension $(100,100,6)$. Each dataset is processed through a dedicated embedding block. For the particle kinematics data, we employ an MLP consisting of three fully connected layers with 256, 128, and 32 neurons, respectively. Each layer is followed by a GELU activation function [62] and Root Mean Square Layer Normalization (RMSNorm) [63]. While different normalization techniques can be used, we found that RMSNorm enhances network convergence. For the pairwise interaction dataset, three two-dimensional convolution layers are employed with 256, 128, and 32 filters. The filter size is 1 such that the mapped features are not diluted by the convolution operation. Each convolution layer is followed by GELU activation and RMSNorm layers. The dimension of the embedded kinematic dataset is $(100,32)$ while the embedded particle interaction dataset has the dimension $(100,100,32)$.
The attention based layer has a distinct structure, processing two input datasets and outputting both datasets scaled by the computed attention. For each dataset, an RMSNorm layer is applied to normalize the input features. The trainable matrices, $W_{1}$ and $W_{2}$, are initialized using a fully connected layer for the particle kinematics dataset and a 2D convolutional layer for the pairwise interaction dataset. Both datasets are scaled using skip connections. For the particle kinematics dataset $X_{i}$, an MLP is added after the attention heads. This MLP consists of an RMSNorm layer followed by a fully connected layer with Sigmoid-Weighted Linear Units (SiLU) activation [64]. The MLP size is set to four times the number of embedded features. For the top tagging and quark-gluon tagging tasks, we use 12 attention layers, each containing 16 attention heads. The output of the final attention layer undergoes an average pooling before the final MLP. The final MLP layer consists of a fully connected layer with 100 neurons and an output layer with one neuron and a sigmoid activation function.
3 Performance of IAFormer
In this section, we validate IAFormer against existing networks for Top tagging and quark-gluon classification. The validation is performed using publicly available datasets, the Top [65] and the quark-gluon dataset [66]. Results for existing networks are taken from their respective references. Additionally, we compare IAFormer performance with two baseline networks, a Plain Transformer and a Transformer with an interaction matrix, both trained from scratch, as described in Appendix A.
3.1 Input variables and data structure
In this analysis, we consider two input datasets, one with the particle momentum features and another one of particle pair interaction features. For the top tagging, up to $200$ constituent particles (hadrons) are retained for each jet, with the 4-momenta $(px,py,pz,E)$ of each particle. From this dataset, we construct the first input dataset with up to $100$ $p_{T}$ ordered jet constituents with eleven features for each particle as:
[…]
$\Delta\eta=\eta-\eta_{\text{jet}}$ : pseudorapidity difference
$\Delta\phi=\phi-\phi_{\text{jet}}$ : azimuthal angle difference
[…]
$\log\left(\frac{p_{T}}{p_{T_{\text{jet}}}}\right)$ : normalized $p_{T}$ (GeV)
$\log\left(\frac{E}{E_{\text{jet}}}\right)$ : normalized energy (GeV)
Additionally, the quark-gluon dataset includes particle ID information, providing six extra features: the charge of each final-state particle and five binary indicators specifying whether a particle is an electron, muon, photon, charged hadron, or neutral hadron. The particle ID is represented using one-hot encoding, where a value of 1 corresponds to the given particle and 0 to all others.
For the pairwise particle interactions, we consider six common features for the top and quark gluon tagging [10]
[…]
Similar to [49], we consider the logarithm of the pairwise interaction variables. Integrating the energy and momentum variables in the interaction matrix is crucial as IAFormer uses the interaction matrix only for the attention score. The first two variables identify the most energetic particle pairs that most likely form the prongs of the Top jet.
The plain Transformer network uses only the particle kinematic dataset with the dimension $(100,11)$, with the first and second numbers referring to the maximum number of jet constituents and the features, respectively. IAFormer uses two input datasets, particle kinematics and pairwise interaction matrix of the dimension of $(100,100,6)$.
3.2 Top tagging
The identification of jets originating from hadronically decaying top quarks, known as top tagging, plays a vital role in LHC new physics searches. To evaluate the performance of our proposed network, we employ the top tagging dataset. This dataset consists of jets generated at a center-of-mass energy of $\sqrt{s}=14$ TeV using Pythia8 [67], with fast detector simulation by Delphes [68]. The simulation does not incorporate effects from multiple parton interactions or pileup. Jet clustering is carried out using the Anti-kt algorithm with a cone radius of $R=0.8$, based on Delphes E-Flow objects. The dataset includes jets with transverse momentum in the range $p_{T}\in[550,650]$ GeV and pseudo-rapidity constrained to $|\eta|<2$. For top quark events, a valid jet must be located within $\Delta R=0.8$ of a hadronically decaying top quark, with all three decay products of the top also confined within $\Delta R=0.8$ from the jet axis. The primary background process considered is QCD dijet production. This dataset consists of one million $t\bar{t}$ events and an equal number of QCD dijet events. Following the standard data split, we allocate 1.2 million events for training, 400,000 for validation, and 400,000 for testing. The dataset is widely used in previous studies, allowing for direct performance comparisons of existing models.  [^2] It should be noted that there is an effective class imbalance near the top quark mass. The $t\bar{t}$ sample exhibits a pronounced peak around 170 GeV, whereas the mass of the QCD jets is concentrated near zero, resulting in tiny overlap between their distributions. This disparity makes it particularly challenging to discern fine-grained differences among high-performance neural networks. Therefore we use several measures to compare the network performance.
Table 1: Performance of different networks for Top jet classification. Results for networks without an asterisk are taken from their respective references. The plain Transformer was trained from scratch using the structure described in Appendix A.
Accuracy AUC $1/\epsilon_{B}(\epsilon_{s}=0.5)$ $1/\epsilon_{B}(\epsilon_{s}=0.3)$ Parameters
ParticleNet[54] 0.940 0.9858 $397\pm 7$ $1615\pm 93$ 370K
PFN[53] -- 0.9819 $247\pm 3$ $888\pm 17$ 86.1K
rPCN[69] -- 0.984 $364\pm 9$ $1642\pm 93$-
Lorentz invariance based networks
PELICAN[35] 0.9426 0.987 -- $2250\pm 75$ 208K
LorentzNet[70] 0.942 0.9868 $498\pm 18$ $2195\pm 173$ 224K
L-GATr[71] 0.942 0.9870 $540\pm 20$ $2240\pm 70$
Attention based networks
ParT[49] 0.940 0.9858 $413\pm 6$ $1602\pm 81$ 2.14M
MIParT[50] 0.942 0.9868 $505\pm 8$ $2010\pm 97$ 720.9K
Mixer[21] 0.940 0.9859 $416\pm 5$ 86.03K
OmniLearn[72] 0.942 0.9872 $568\pm 9$ $2647\pm 192$ 1.6M
Plain Transformer 0.927 0.979 $362\pm 7$ $780\pm 73$ 1.7M
IAFormer 0.942 0.987 $510\pm 6$ $2012\pm 30$ 211K
Table 1 presents the classification performance of IAFormer for top jet classification. IAFormer is trained for 23 epochs, with a batch size of 256 with early stopping set for 5 epochs. The AdamW optimizer [73] with an initial learning rate of $5\times 10^{-4}$ and the Cosine annealing scheduler is used to adjust the learning rate during the training. Two input datasets are considered, particle kinematics, with dimensions $X=(100,11)$, with one-hundred maximum number of particles and 11 features per particle. A second dataset of the pairwise interaction $\cal{I}$, with dimension $(100,100,6)$, which encodes 6 features for each particle pair. Other networks are trained with the same hyperparameter setup, but trained for 20 epochs. After training, the network performance is evaluated using various metrics, including classification accuracy, Area Under the ROC Curve (AUC), and background rejection at signal efficiencies of 0.3 and 0.5.
IAFormer outperforms other attention-based Transformer networks, although it has an order of magnitude smaller parameter size of 211K than ParT. This reduction in network size is achieved by replacing the $Q$ and $K$ matrices with an interaction matrix, where attention scores are primarily based on pairwise particle interactions. Furthermore, the use of sparse attention enables the network to suppress attention scores of less relevant tokens, reducing the need for excessive model complexity, while efficiently distinguishing between Top and QCD jets.
Figure 2: Left plot shows the $\beta$ distribution across the twelve IAFormer layers for different seed numbers on the test dataset. Right plot illustrates the uncertainty in the network output when trained with different seed numbers, 42, 0, 7, 123, and 12345. The architectures of the Transformer $+I_{ij}$ and the plain Transformer. The uncertainty is reported in terms of background rejection at 0.3 signal efficiency.
This is reflected in the network training uncertainty in which IAFormer is much robust against numerical fluctuation. The right plot of figure 2 illustrates the fluctuation of the background rejection at 0.3 signal efficiency with the different seeds for the network parameter initialization. Furthermore, IAFormer exhibits robustness against the change of the seed number with a fluctuation range of $150$ while plain Transformer has a fluctuation range of $600$ and Transformer with interaction matrix has a fluctuation range of $550$.
The core component of the dynamic sparse attention is the inclusion of the learnable parameter $\beta$, which regulates the level of suppression of less relevant tokens. The left plot in Figure 2 illustrates the distribution of $\beta$ across all IAFormer layers. To better understand the role of $\beta$, we analyze three different random seeds. Interestingly, all distributions exhibit a similar pattern: $\beta$ values increase in the initial layers, reach a maximum value and then decrease in the later layers. Network classification accuracy improves for higher $\beta$; the blue line corresponds to the best classification accuracy of AUC=0.98678, with $\beta$ value very close to 0.6. It also starts at a lower value for earlier layers. This shows the earlier layer tries to build collective quantities stable against the fluctuations, and the rest of the information is abandoned in the later layers dynamically for successful training.
3.3 Quark gluon tagging
The quark jet and gluon jet are expected to show distinguishable features due to different fragmentation patterns, due to the color difference. Because BSMs tend to emit high-energy quarks rather than gluons, the quark-gluon tagging has been considered to be phenomenologically important. The public dataset [66] is generated using Pythia8 [67], where quark jets originate from $Z^{0}$ boson production plus quark processes, while gluon jets stem from $Z^{0}$ boson production plus gluon events. The dataset contains not only the four-momentum information but also particle id (PID), which can significantly enhance jet identification performance. The PID information is incorporated in a way that aligns with the experimental setup, categorizing particles into five broad types: electrons, muons, charged hadrons, neutral hadrons, and photons, along with their electric charge. These six PID attributes are incorporated as one hot vector, and they are combined with the seven kinematic variables, resulting in a total dimension of 13 input features per particle. These final state particles are clustered into jets using the anti-$k_{T}$ algorithm with a radius parameter of $R=0.4$. Only jets with transverse momentum $p_{T}$ in the range [500, 550] GeV and rapidity $|y|<2$ are included. The dataset consists of 2 million jets, evenly split between signal and background. Following the recommended partitioning, we allocate 1.6 million jets for training, 200,000 for validation, and 200,000 for testing for IAFormer and plain Transformer.
Table 2: Performance of different networks on the quark-gluon classification task. Results for networks without an asterisk are taken from their respective references. The plain Transformer was trained from scratch using the architecture described in Appendix A. The results for ParT and MIParT are shown in italic letters, because they include additional scaling factors for charged and neutral hadrons[49], which are not used in the other networks.
Accuracy AUC $1/\epsilon_{B}(\epsilon_{s}=0.5)$ $1/\epsilon_{B}(\epsilon_{s}=0.3)$ Parameters
ParticleNet[54] 0.840 0.9116 $39.8\pm 0.2$ $98.6\pm 1.3$ 370K
PFN[53] -- 0.9052 $37.4\pm 0.7$- 86.1K
rPCN[69] -- 0.9081 $38.6\pm 0.5$- --
Lorentz invariance based networks
PELICAN[35] 0.8551 0.9252 ${52.3\pm 0.3}$ ${149.8\pm 2.4}$ 208K
LorentzNet[70] 0.844 0.9156 $42.4\pm 0.4$ $110.2\pm 1.3$ 224K
Attention based networks
ParT[49] 0.849 0.9203 $\mathit{47.9\pm 0.5}$ $\mathit{129.5\pm 0.9}$ 2.14M
MIParT[50] 0.851 0.9215 $\mathit{49.3\pm 0.4}$ $\mathit{133.9\pm 1.4}$ 720.9K
OmniLearn[72] 0.844 0.9159 $43.7\pm 0.3$ $107.7\pm 1.5$ 1.6M
Plain Transformer 0.831 0.890 $32.6\pm 0.5$ $86.\pm 1.7$ 1.7M
IAFormer 0.844 0.917 $42\pm 0.3$ $101\pm 1.1$ 171K
[…]
Figure 3: $\beta$ distribution for quark-gluon test data set for three different seed number.
IAFormer architecture was optimized for best performance using 6 attention layers, with a total parameter count of 171k. The reduction in network size, compared to the top tagging case, is due to the similarity between quark and gluon jets. In this scenario, the network performance saturates after a certain number of layers, and adding more layers only introduces redundancy without improving accuracy. The optimization of the network size is based on the distribution of $\beta$ parameters for the attention layers. For IAFormer with 12 layers, $\beta$ distribution randomly fluctuates across layers with little improvement in the classification accuracy. By reducing the network size, the $\beta$ parameter shows a similar distribution as the Top tagging case, as shown in Figure 3.
Finally, the obtained $\beta$ value of the last layer is larger than the top vs QCD classification of the previous subsections. Because $\beta$ appears negatively in the attentions, the larger $\beta$ likely indicates that the effective degree required for the quark-gluon classification is smaller than the top-QCD classification.
4 Analysis of hidden layer representations
We now compare the attention-based transformer models (Plain Transformer, Transformer + $\mathcal{I}_{ij}$ and IAformer) using several interpretability methods to study how hidden layers are organized for classification. Different interpretability methods can be applied to analyze attention-based Transformer models designed for particle cloud data. These techniques provide insights into how the model processes and prioritizes different particles, helping to understand learned representations and the network decision-making. By leveraging these interpretability techniques, one can gain deeper insights into the behavior of attention-based Transformers for particle cloud analysis, ensuring that their predictions align with meaningful physical patterns. Common interpretation methods are
  • Attention Maps: These maps visualize how the Transformer model distributes its attention across different particles in the cloud [74]. By highlighting the most influential particles, attention maps provide an intuitive understanding of the pair of tokens relevant to the classifications.
  • CKA Similarity: Centered Kernel Alignment (CKA) measures the similarity between hidden layer representations within a model or across different models [75]. A high CKA similarity between layers indicates redundancy, suggesting that certain layers can be pruned without significant loss of performance, while a low CKA similarity suggests that layers learn distinct features that contribute to improved classification accuracy.
  • Saliency Maps: These maps measure the sensitivity of the model’s output to small perturbations in the input [76]. Saliency maps help assess whether the model captures physically meaningful features or relies on spurious correlations.
  • Layer-wise Relevance Propagation (LRP): LRP assigns relevance scores to individual particles, quantifying their contribution to the final decision [77]. This technique helps trace back model predictions to specific input particles. Different LRP variants can be used to enhance interpretability for high-energy physics applications.
  • Grad-CAM: This method generates class-specific activation maps by computing the gradients of the predicted class score with respect to the final Transformer layer [78]. For example, Grad-CAM can highlight crucial regions in the $eta-\phi$ plane, allowing for a geometric interpretation of how the model distinguishes between different classes.
In the following, we focus on attention maps and CKA similarity to interpret the results obtained from the Top tagging. The attention maps utilized to visualize the sparse pattern in the attention score when using IAFormer, while the CKA is used to explore the learned representation by the different attention layers.
4.1 Attention Maps
The attention map is a 2D heatmap of attention scores that highlights where the model focuses. To compute attention maps, the attention scores, which represent the strength of the relationship between different elements of the input, are computed by taking the dot product between the query vector of one element and the key vector of every other element in the input. These scores are then normalized using a softmax function, which ensures that they sum to one. The final attention map is a visual representation of these scores, showing which parts of the input the model “attends to” or focuses on most when producing its output. In IAFormer the attention map is computed as the difference between the learnable pairwise interaction matrix in equation 9.
Figure 4: Attention maps of the final self-attention layer of IAFormer (left) and a plain Transformer (right), generated using 10,000 test events from the Top jet dataset. Each network comprises 16 attention heads, shown individually. The axes represent particle tokens, with each event containing 100 particles, while the color bar denotes the attention scores assigned to each particle pair, ranging from 0 to 1.
Figure 4 presents the attention maps from the final layer of IAFormer and a plain Transformer, using 10,000 test events. In both networks, the final layer consists of 16 attention heads, displayed individually. In these maps, the $X$ and $Y$ axes represent the particle tokens, hadrons within the jet, while the color scale indicates the attention score assigned to each particle pair. It should be noted that the tokens do not exactly correspond to the particle inputs, because the token of the final layer is a mixture of input particles.
The attention maps for IAFormer show high attention scores concentrated on a few particle tokens across all attention heads, with the remaining particles significantly suppressed. In contrast, the plain Transformer distributes attention more broadly across a larger number of tokens. This shows IAFormers more efficiently accumulate relevant information from earlier tokens. Ideally, the network should focus on the cluster hadrons that form the three-prong structures of the top jet, as assigning high attention to a large number of tokens can hinder classification performance. Additionally, IAFormer attention maps exhibit a symmetric pattern, reflecting the symmetric nature of the pairwise interaction matrix. This symmetry arises even though we do not require the symmetry to weight factor $W_{1}$ and $W_{2}$.
4.2 CKA similarity
Centered Kernel Alignment (CKA) similarity, derived from kernel methods and alignment-based metrics, provides a powerful tool for assessing how closely two sets of learned representations correspond, whether across different layers within a neural network or between entirely separate models. Unlike traditional similarity measures such as Pearson correlation or Euclidean distance, which directly compare raw values, CKA evaluates how well representations align in high-dimensional feature space. This allows it to capture intricate, nonlinear relationships between features, making it particularly effective for analyzing deep learning models. One of the key challenges in studying neural network representations is that features are distributed across multiple neurons, and the number of neurons in hidden layers often exceeds the input dimension while varying across layers and models. This variability makes direct comparisons difficult. CKA addresses this issue by providing a quantitative framework for comparing learned representations, both within a single model and across different architectures.
To compute CKA similarity, consider two activation matrices, $X\in\mathbb{R}^{d\times P_{1}}$ and $Y\in\mathbb{R}^{d\times P_{2}}$, where $d$ represents the number of input samples, and $P_{1}$ and $P_{2}$ denote the number of neurons in two different hidden layers. The CKA similarity is defined as
$\text{CKA}(M,N)=\frac{\text{HSIC}(M,N)}{\sqrt{\text{HSIC}(M,M)\text{HSIC}(N,N)}}\,,$ (12)
with $M=XX^{T}$ and $N=YY^{T}$ are the Gram matrices of the two hidden layers, both of size $d\times d$. Since the Gram matrix dimensions depend only on the number of input samples, CKA allows for the comparison of hidden layers with different neuron counts, as well as representations from different models. The Hilbert-Schmidt Independence Criterion (HSIC) measures statistical dependence between two matrices and is defined as
$\text{HSIC}(M,N)=\frac{1}{(d-1)^{2}}\text{Tr}(MHNH)\,,$ (13)
where $H_{ij}=\delta_{ij}-\frac{1}{d}$ is a centering matrix, which ensures that each row and column sums to zero for $A=M,N$. Centering prevents CKA from being biased by extreme values or outliers, ensuring more stable and meaningful comparisons.
CKA values range from 00 to $1$, where higher values indicate strong similarity between learned representations. If two consecutive layers have high CKA similarity, it suggests that the second layer does not significantly enhance classification accuracy, implying that it could be removed without affecting performance. On the other hand, layers with low CKA similarity capture distinct aspects of the data, contributing to improved model performance by learning complementary features.
Figure 5: Linear CKA similarity for IAFormer (left), Transformer $+I_{i,j}$ (middle) and Plain Transformer (right) using 1000 test events from the Top jet dataset. The axes represent the attention layers in each network, while the colour bar indicates the CKA values.
We compute the output from each attention layer with dimensions $(100,32)$, where the first dimension corresponds to the number of particles and the second represents feature embeddings. To construct the Gram matrices, $M$ and $N$, we consider 1000 test events and average over the feature dimension, resulting in Gram matrices of size $(1000,1000)$. These matrices are then used to compute the CKA values for three different networks, IAFormer, Transformer with an interaction matrix, and a plain Transformer, as shown in Figure 5. For a fair comparison, all networks are trained with 12 attention layers, represented on the $X$ and $Y$ axes of each plot.
Attention layers of IAFormer exhibit lower CKA values, particularly in the first four layers, with CKA values ranging from 0.5 to 0.8. This suggests that these layers capture different patterns in the jet constituents, contributing to an overall improvement in classification performance. The later layers exhibit stronger internal similarity, with CKA values exceeding 0.8. The CKA structure indicates an efficient flow of attention from the early layers to the final ones. The middle plot presents the CKA values for the Transformer incorporating the interaction matrix. The CKA matrix exhibits a block-diagonal structure, where the first four layers demonstrate strong internal similarity but differ significantly from the remaining layers. Additionally, the last four layers encode nearly identical information, with a CKA value of 1. In contrast, the plain Transformer shows consistently high similarity across all layers, with CKA values ranging from 0.85 to 1. This uniformity likely contributes to its lower performance in the Top tagging task compared to other networks.
Overall, the Plain Transformer and Transformer+ $\mathcal{I}_{ij}$ stop improving first 6 layers, while IAFormer steadily improves for all layers, with stronger suppression of the fluctuation by larger $\beta$ layer by layer, likely building global variables along the increasing layers.
5 Conclusion
In this paper, we introduce IAFormer, an attention-based deep learning network that incorporates dynamic sparse attention to reduce network size while preserving classification performance. Furthermore, the standard attention weights are replaced with pairwise interactions between input particles. Combined with sparse attention, this modification reduces redundancy typically present in standard Transformer architectures.
To evaluate the network performance, we validate IAFormer on two benchmark tasks: top quark tagging and quark–gluon discrimination. In both cases, IAFormer achieves improved classification performance compared to the ParT network, while significantly reducing model size. Furthermore, IAFormer delivers performance competitive with state-of-the-art Lorentz-invariant networks. To further analyze the results, we employ various AI interpretability techniques to visualize the information learned by IAFormer’s hidden layers in comparison to other Transformer-based architectures. Specifically, we utilize attention maps and CKA similarity. The attention maps reveal a sparse structure in IAFormer attention weights, in contrast to the more scattered attention patterns observed in the original Transformer, where weights are more uniformly distributed across all particle tokens. This sparsity enables IAFormer to concentrate on the most relevant particles and suppress attention to less informative inputs, a behavior not exhibited by the original Transformer, which tends to assign attention broadly, even to less important tokens. Additionally, CKA analysis indicates lower similarity between IAFormer learned representations and those of other Transformer variants, highlighting its unique learning dynamics.
Although IAFormer demonstrates improved performance over existing Transformer-based networks, its architecture still requires optimization for specific tasks. For instance, the learnable parameter $\beta$ can exceed 1, which amplifies noise in the attention weights. In IAFormer, the value of $\beta$ is clipped to a maximum of 1 to mitigate this effect. Additionally, it is possible for attention scores to become negative, making them unsuitable for interpretation as attention probabilities.
In fact, it may be worthwhile to consider the physical interpretation of the trainable parameter $\beta_{i}$. These hyperparameters depend on the underlying probability distributions of the signal and background, and can be viewed as global variables that quantify the distinction between different jet classes. In particular, the values of $\beta$ in the final layers are likely associated with the effective degrees of freedom necessary to capture the differences between signal and background, assuming the network depth has been appropriately optimized.
Finally, various IAFormer architectures tailored for different classification tasks have been developed and made publicly available.
Acknowledgments
This work is funded by grant number 22H05113, “Foundation of Machine Learning Physics”, Grant in Aid for Transformative Research Areas and 22K03626, Grant-in-Aid for Scientific Research (C). WE is funded by the ErUM-WAVE project 05D2022 “ErUM-Wave: Antizipation 3-dimensionaler Wellenfelder”, which is supported by the German Federal Ministry of Education and Research (BMBF).
Appendix A Public code
Although IAFormer is validated against the current DL networks for jet tagging problems, it is adopted for generic classification problems. Additionally, our code comprises three different networks that can be used by public users, IAFormer, Plain Transformer and Transformer with interaction matrix. In this section, we present the full details on how to use the implemented networks for different classification tasks. Implementation of the IAFormer model is publicly available at IAFormer. The repository is implemented using PyTorch Lightning [^3] https://github.com/Lightning-AI/pytorch-lightning, a high-level deep learning framework built on top of PyTorch [79]. It offers a modular, scalable, and user-friendly interface for training and evaluating deep learning models.
The repository is organized into several key components:
  • utils/graph_builder.py: A module for transforming raw jet constituent data into an input suitable format for the IAFormer.
  • models/: Contains the implementation of the IAFormer model, including the implementation of vanilla multi-head attention and sparse attention.
  • data/: Implementation of data loading and preprocessing.
  • configs/: YAML configuration files for different transformer based models.
  • callbacks/: Implementation of inference and plotting.
Installation instructions and guidelines for training and evaluating different models are provided in the GitHub repository’s README.md. Users can easily reproduce results or experiment with custom configurations by modifying the corresponding YAML files. Training the network is done via the following command
python main.py fit -config=configs/ia_former.yaml
The code is equipped with three different configuration files for the different networks. Hyperparamters of the best performance are saved in the checkpoints directory. To test the network
python main.py test -config=configs/ia_former.yaml -ckpt_path=best_point.ckpt
In the following, we describe the individual components and modules that constitute the IAFormer implementation.
A.1 Particle Graph Construction
We have implemented two dedicated preprocessing classes: TopParticleGraphBuilder and QGParticleGraphBuilder for the two different datasets used in this article, the top tagging dataset and the quark-gluon dataset. Each class performs the same task, but is tailored to the input format of its respective dataset.
Each class reads the four-momentum components (PX, PY, PZ, E) of the jet constituents for each event, computes derived kinematic quantities and pairwise interaction features, and stores the output in an HDF5 file, which contains the following datasets:
  • feature_matrix: A 3D array of shape $[N_{\text{events}},N_{\text{particles}},F]$. Here, $F$ are the features, which incldues the particle four momenta, log-scaled transverse momentum and energy, relative kinematics, spatial distances, (Eq. 10) and particle identification variables, for the quark-gluon dataset.
  • adjacancy_matrix: A 3D array of shape $[N_{\text{events}},(N_{\text{particles}}\cdot(N_{\text{particles}}-1))/2,6]$ that contains pairwise interaction features between particles, which are listed in Eq. 11.
  • mask: A 2D boolean array of shape $[N_{\text{events}},N_{\text{particles}}]$ indicating valid (non-padded) particle entries in each event.
  • labels: A 1D array of shape $[N_{\text{events}}]$ that stores the ground-truth class labels.
A.2 Data Handling Module
The data_handling.py module provides functionality for loading the preprocessed HDF5 files generated by the graph_builder.py module and integrating them with PyTorch Lightning workflow.The data loading class H5Dataset implements the reconstruct_adjacency() method, which infers the number of particles from the number of pairwise entries and reconstructs the dense adjacency matrix accordingly.
A.3 Networks structure
The core classification model used in IAFormer is implemented in the Transformer. It supports multiple attention mechanisms mentioned above via the attn argument. The training logic is handled by the JetTaggingModule, a PyTorch Lightning module that encapsulates the training, validation, and testing procedures.
Transformer Class
The Transformer model consists of the following components:
  • Input Embeddings:
    • ParticleEmbedding: Projects node-level features into a common embedding space using an MLP.
    • InteractionInputEncoding: Projects edge-level features ($I_{i,j}$) into a common embedding space using a series of Conv2D layers.
  • Attention Blocks: A stack of ParticleAttentionBlock modules.
  • Feature Aggregation and Output: The per-particle representations are pooled across the feature dimension, and passed to a linear classification head for binary classification.
JetTaggingModule
The JetTaggingModule is a PyTorch Lightning wrapper around the Transformer model. It defines the full training and evaluation pipeline:
  • training_step(): Computes binary cross-entropy loss on the logits and logs training accuracy.
  • validation_step() and test_step(): Evaluate the model on validation and test datasets, respectively, computing accuracy, loss and the area under the ROC curve.
During testing, attention weights and intermediate activations from each transformer block are collected for later downstream analysis. The following arguments control the architecture and training process:
  • embed_dim: Dimension of the embedding space for particles.
  • num_heads: Number of attention heads per block.
  • num_blocks: Depth of the transformer.
  • max_num_particles: Maximum number of particles per event.
  • attn: Type of attention used (plain, interaction, or sparse).
A.4 Attention Module
The attention.py module implements three types of attention mechanisms used in this article:
  • Plain Attention: A standard scaled dot-product attention without any interaction features.
  • Interaction Attention: Extends plain attention by incorporating an interaction matrix into the attention logits before the softmax operation.
  • Sparse Differential Attention: A novel attention mechanism that combines two interaction-aware attention maps using a learnable $\beta$.
MultiHeadAttention
The MultiHeadAttention class implements standard multi-head attention. It supports an optional interaction matrix $I_{i,j}$ that is projected using a Conv2D layer to match the shape of the attention logits. If provided, this interaction matrix is added to the dot-product attention scores before applying softmax. An optional attention mask (umask) can be used to ignore invalid or padded elements.
The module consists of:
  • Linear projections for queries, keys, and values: q_proj, k_proj, v_proj.
  • A learned projection for the interaction matrix: u_proj.
  • A final output projection: out_proj.
MultiHeadSparseAttention
The MultiHeadSparseAttention class implements a differential attention mechanism. The attention map is calculated entirely from the projected interaction matrix $I_{i,j}$ using a Conv2D layer. Two attention maps are extracted and combined via a learnable scalar $\beta$, parameterized through dot products between $\beta_{q}$ and $\beta_{k}$. This allows the model to softly subtract one attention map from another, enabling a sparse or selective focus on particle-particle interactions.
  • The learnable $\beta$ coefficient is initialized using a depth-dependent function: beta_init.
  • The attention weights are computed as:
    $\text{attn}=\text{Softmax}(I_{i,j})[:,:,0]-\beta\cdot\text{Softmax}(I_{i,j})[:,:,1]$
ParticleAttentionBlock
The ParticleAttentionBlock class is a modular block that wraps the attention mechanism with normalization and a feed-forward MLP:
  • It supports all three attention variants via the attn argument (plain, interaction, diff).
  • Applies residual connections and normalization both before and after the attention sub-layer.
  • Includes an MLP with GELU activation and expansion factor.